Circular No. 8414

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004es AND 2004et A. V. Filippenko and R. J. Foley, University of California at Berkeley

(UCB), report that inspection of CCD spectra (range 330–1000 nm), obtained by T. Treu and M. A. Malkan on Oct. 1 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2004es (IAUC 8412) is of type II, perhaps 1–2 months past maximum brightness. The Hα line has little or no absorption component, but the emission component is strong and broad. Na I D and some Fe II lines have P-Cyg profiles. SN 2004et (IAUC 8413) is also of type II, confirming the classification derived from the high-resolution spectrum reported on IAUC 8413. The P-Cyg profile of Hα is greatly dominated by the emission component, but the other hydrogen Balmer lines have more typical P-Cyg profiles. The continuum is quite blue, although it plummets shortward of 400 nm. Narrow interstellar Na I D and Ca II H and K absorption lines are prominent.

W. Li and A. V. Filippenko, UCB; S. D. Van Dyk, Spitzer Science Center, California Institute of Technology; and J.-C. Cuillandre, Canada-France-Hawaii Telescope (CFHT) Corporation, report that their analysis of deep, high-quality (seeing 0′′.8) images of NGC 6946, taken with the 3.6m CFHT on 2002 Aug. 6 UT, shows a possible progenitor for SN 2004et consistent in position with the more blurred, blended object previously found in lower-quality images (seeing 2′′.7; IAUC 8413) — located at α = 20^{h}35^{m}25^{s}.41, δ = +60^{o}07′18′′.1 (equinox 2000.0; uncertainty ∼ 0′′.2 in each coordinate), within 0′′.3 of the position of SN 2004et measured from KAIT images (IAUC 8413), with magnitudes B = 23.84 ± 0.22, V = 22.83 ± 0.26, and R = 22.31 ± 0.18. Using the same distance and reddening estimates to SN 2004et (cf. IAUC 8413), the possible progenitor has M(B) = −6.5, M(V) = −7.1, and M(R) = −7.4, which may be consistent with a massive supergiant, although likely too bright and too blue for a single red supergiant. However, with ground-based data it is difficult or impossible to determine accurately the exact properties of the progenitor, given its probable membership in a compact association. Note that both the distance and reddening to SN 2004et may have substantial uncertainties.

COMET 160P/2004 NL_{21} (LINEAR) Following the identification by M. Meyer (cf. MPEC 2004-S18) of 1996

observations of comet P/2004 NL_{21} (cf. IAUC 8408), on exposures taken via NEAT at Haleakala, the comet has received the permanent number 160P (cf. MPC 52734, 52764, 52766).

2004 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green